# Proximal Oracles for Optimization and Sampling

### v1

@@ 5 Proximal sampling algorithm
As discussed earlier, to use ASF in sampling problems, we need to realize the RGO with efficient implementations. In the rest of this section, we develop efficient algorithms for RGO associated with the two scenarios of sampling we are interested in, and then combine them with the ASF to establish a proximal algorithm for sampling. The complexity of the proximal algorithm can be obtained by combining the above convergence results for ASF and the complexity results we develop for RGO. The rest of the section is organized as follows. In Subsection 5.1 we develop an efficient algorithm for RGO associated with semi-smooth potentials via rejection sampling. This is combined with ASF to obtain an efficient sampling algorithm from semi-smooth potentials. In Subsection 5.2, we further extend results to the second setting, i.e., composite potentials.
5.1 Sampling from semi-smooth potentials
The bottleneck of using the ASF (Algorithm 1) in sampling tasks with general distributions is the availability of RGO implementations. In this subsection, we address this issue for convex semi-smooth potentials by developing an efficient algorithm for the corresponding RGO.
@@ 5.1 Sampling from semi-smooth potentials
  4. If
$U\leq\frac{\exp(-f_{y}^{\eta}(X))}{\exp(-h_{1}(X))},$
then accept/return $X$; otherwise, reject $X$ and go to step 2.
@@
Assume $f$ is convex and $L_{\alpha}$-semi-smooth. Let $f_{y}^{\eta}$ be as in (9) and $h_{1}$ and $h_{2}$ be as in (32). Then, for every $x\in\mathbb{R}^{d}$, we have
@@
Proof: The first inequality in (33) immediately follows from Lemma 3.1(d) and the definition of $h_{1}$ in (32a). By the definition of $f_{y}^{\eta}$ in (9) we get
@@
Using the above inequality and the definition of $h_{2}$ in (32b), we conclude that the second inequality in (33) holds.
@@
Proof: Denote $r=\|x\|$, then
@@
The result (36) then follows from the above inequality and (37).
@@
Assume $f$ is convex and $L_{\alpha}$-semi-smooth. If
@@
Proof: It is a well-known result for rejection sampling that $X\sim\pi^{X|Y}(x\mid y)$ and the probability that $X$ is accepted is
@@
We finally bound the total complexity to sample from a log-concave distribution $\nu$ in (2) with a semi-smooth potential $f$. We combine our efficient algorithm (Algorithm 4) of RGO for semi-smooth potentials and the convergent results for ASF, namely Theorem 5.1, to achieve this goal.
@@
Assume $f$ is convex and $L_{\alpha}$-semi-smooth, then Algorithm 1, initialized with $\rho_{0}^{X}$ and stepsize $\eta\asymp 1/(L_{\alpha}^{\frac{2}{\alpha+1}}d)$, using Algorithm 4 as an RGO has the iteration-complexity bound
@@
Proof: The results follow directly from Theorem 5.1, Theorem 3.5 and Proposition 5.4 with the choice of stepsize $\eta\asymp 1/(L_{\alpha}^{\frac{2}{\alpha+1}}d)$.
5.2 Sampling from composite potentials
In this subsection, we consider sampling from a log-concave distribution $\nu\propto\exp(-f(x))$ associated with a composite potential $f$ satisfying (6). This setting is a generalization of the semi-smooth setting studied in the previous sections. It turns out that both Algorithm 1 and the implementation for RGO, Algorithm 4, developed for semi-smooth sampling can be applied directly to this general setting with properly chosen stepsizes. Below, we extend the analysis in Subsection 5.1 to the composite setting and establish corresponding complexity results.
The following lemma is a counterpart of Lemma 5.2 in the composite setting. Its proof is given in Appendix B.
@@ 5.2 Sampling from composite potentials
Proof: Using the Young’s inequality $st\leq s^{p}/p+t^{q}/q$ with
@@
Through the above arguments, we show that Algorithm 4 designed for semi-smooth potentials is equally effective for composite potentials satisfying (6). Combining Proposition 5.8 and Theorem 3.9 with the convergence results for ASF, we obtain the following iteration-complexity bounds for sampling from composite potentials. The proof is similar to that of Theorem 5.5 and is thus omitted.
@@
Assume $f$ is a convex and satisfies (6). Consider Algorithm 1, initialized with $\rho_{0}^{X}$ and stepsize $\eta$ satisfies (52), using Algorithm 4 as a RGO. Each RGO requires $\tilde{\cal O}\left(\frac{1}{d\delta}+1\right)$ subgradient evaluations of $f$ and $\exp\left(\delta+\frac{1}{2}+\frac{\sum_{i=1}^{n}(1-\alpha_{i})}{4}\right)$ rejection steps in expectation. The total complexity of Algorithm 1 to achieve $\varepsilon$ error in terms of KL divergence is
@@
6 Conclusion
In this paper, we study proximal algorithms for both optimization and sampling lacking smoothness. We first establish the complexity bounds of the regularized cutting-plane method for solving proximal subproblem (3), where $f$ is convex and satisfies either (5) (semi-smooth) or (6) (composite). This efficient implementation gives an approximate solution to the proximal map in optimization, which is the core of both proximal optimization and sampling algorithms.
@@ 6 Conclusion
For sampling, we propose an efficient method based on rejection sampling and the approximate proximal map to realize the RGO, which is a proximal sampling oracle. Finally, combining the sampling complexity of RGO and the complexity bounds of ASF, which is a counterpart of the proximal point method in sampling, we establish the complexity bounds of the proximal sampling algorithm in both semi-smooth and composite settings.

### v2

@@ 5 Proximal sampling algorithm
As discussed earlier, to use ASF in sampling problems, we need to realize the RGO with efficient implementations. In the rest of this section, we develop efficient algorithms for RGO associated with the two scenarios of sampling we are interested in, and then combine them with the ASF to establish a proximal algorithm for sampling. The complexity of the proximal algorithm can be obtained by combining the above convergence results for ASF and the complexity results we develop for RGO. The rest of the section is organized as follows. In Subsection 5.1 we develop an efficient algorithm for RGO associated with Hölder smooth potentials via rejection sampling. This is combined with ASF to obtain an efficient sampling algorithm from Hölder smooth potentials. In Subsection 5.2, we further extend results to the second setting, i.e., hybrid potentials.
5.1 Sampling from Hölder smooth potentials
The bottleneck of using the ASF (Algorithm 1) in sampling tasks with general distributions is the availability of RGO implementations. In this subsection, we address this issue for convex Hölder smooth potentials by developing an efficient algorithm for the corresponding RGO.
@@ 5.1 Sampling from Hölder smooth potentials
  if $U\leq\exp\left(-f_{y}^{\eta}(X)+h_{1}(X)\right)$then
     accept/return $X$;
  else
     reject $X$ and go to step 2.
  end if
@@
Assume $f$ is convex and $L_{\alpha}$-Hölder smooth. Let $f_{y}^{\eta}$ be as in (9) and $h_{1}$ and $h_{2}$ be as in (33). Then, for every $x\in\mathbb{R}^{d}$, we have
@@
Proof The first inequality in (34) immediately follows from Lemma 3.1(d) and the definition of $h_{1}$ in (33a). By the definition of $f_{y}^{\eta}$ in (9) we get
@@
Using the above inequality and the definition of $h_{2}$ in (33b), we conclude that the second inequality in (34) holds. ∎
@@
Proof Denote $r=\|x\|$, then
@@
The result (37) then follows from the above inequality and (38). ∎
@@
Assume $f$ is convex and $L_{\alpha}$-Hölder smooth. If
@@
Proof It is a well-known result for rejection sampling that $X\sim\pi^{X|Y}(x\mid y)$ and the probability that $X$ is accepted is
@@
We finally bound the total complexity to sample from a log-concave distribution $\nu$ in (2) with a Hölder smooth potential $f$. We combine our efficient algorithm (Algorithm 4) of RGO for Hölder smooth potentials and the convergent results for ASF, namely Theorem 5.1, to achieve this goal.
@@
Assume $f$ is convex and $L_{\alpha}$-Hölder smooth, then Algorithm 1, initialized with $\rho_{0}^{X}$ and stepsize $\eta\asymp 1/(L_{\alpha}^{\frac{2}{\alpha+1}}d)$, using Algorithm 4 as an RGO has the iteration-complexity bound
@@
Proof The results follow directly from Theorem 5.1, Theorem 3.5 and Proposition 5.4 with the choice of stepsize $\eta\asymp 1/(L_{\alpha}^{\frac{2}{\alpha+1}}d)$. ∎
5.2 Sampling from hybrid potentials
In this subsection, we consider sampling from a log-concave distribution $\nu\propto\exp(-f(x))$ associated with a hybrid potential $f$ satisfying (6). This setting is a generalization of the Hölder smooth setting studied in the previous sections. It turns out that both Algorithm 1 and the implementation for RGO, Algorithm 4, developed for Hölder smooth sampling can be applied directly to this general setting with properly chosen stepsizes. Below, we extend the analysis in Subsection 5.1 to the hybrid setting and establish corresponding complexity results.
The following lemma is a counterpart of Lemma 5.2 in the hybrid setting. Its proof is given in Appendix B.
@@ 5.2 Sampling from hybrid potentials
Proof Using the Young’s inequality $st\leq s^{p}/p+t^{q}/q$ with
@@
Through the above arguments, we show that Algorithm 4 designed for Hölder smooth potentials is equally effective for hybrid potentials satisfying (6). Combining Proposition 5.8 and Theorem 3.9 with the convergence results for ASF, we obtain the following iteration-complexity bounds for sampling from hybrid potentials. The proof is similar to that of Theorem 5.5 and is thus omitted.
@@
Assume $f$ is a convex and satisfies (6). Consider Algorithm 1, initialized with $\rho_{0}^{X}$ and stepsize $\eta$ satisfies (53), using Algorithm 4 as a RGO. Each RGO requires $\tilde{\cal O}\left(\frac{1}{d\delta}+1\right)$ subgradient evaluations of $f$ and in expectation $\exp\left(\delta+\frac{1}{2}+\frac{\sum_{i=1}^{n}(1-\alpha_{i})}{4}\right)$ rejection steps. The total complexity of Algorithm 1 to achieve $\varepsilon$ error in terms of KL divergence is
@@
6 Conclusions
In this paper, we study proximal algorithms for both optimization and sampling lacking smoothness. We first establish the complexity bounds of the regularized cutting-plane method for solving proximal subproblem (3), where $f$ is convex and satisfies either (5) (Hölder smooth) or (6) (hybrid). This efficient implementation gives an approximate solution to the proximal map in optimization, which is the core of both proximal optimization and sampling algorithms.
@@ 6 Conclusions
For sampling, we propose an efficient method based on rejection sampling and the approximate proximal map to realize the RGO, which is a proximal sampling oracle. Finally, combining the sampling complexity of RGO and the complexity bounds of ASF, which is a counterpart of the proximal point method in sampling, we establish the complexity bounds of the proximal sampling algorithm in both Hölder smooth and hybrid settings.